A committee of 4 is selected from a group of 9 faculty and 45 students. In how many different ways can the committee be selected if it must consist of one faculty and three students?
Answer #1

A committee of 4 is selected from a group of 9 faculty and 45 students.

Therefore, 1 faculty can be selected from 9 faculty by,

9C1= 9 ways

And 3 students can be selected from 45 students by,

45C3 = 14190 ways

Therefore, number of different ways can the comittee be selected if it must consist of 1 faculty and 3 students are,

9C1 * 45C3 = 9 * 14190 = 127710

Answer: 127710